In the normal conducting linac, beam commissioning activities continue according to schedule. Plasma conditioning in the ion source and reconditioning of the RFQ section have started. Beam extraction is next.
Linac Warm Units
Particle-free installation of Linac Warm Units is under way in the ESS Accelerator tunnel, with the works performed by the ESS Vacuum team inside a clean room confinement, using a dummy pipe representing the cryomodules. In total, there will be 43 superconducting and 28 normal conducting Linac Warm Units. Set-up of the clean room is prepared every Friday afternoon for particles to settle over the weekend and installation to be performed on Mondays.
RF power and cooling tests of the second serial-produced medium beta cryomodule is ongoing, with 2K helium flow to the cavities of a cryomodule inside a test stand bunker.
NN Bar and port tubes
The so called NN bar, a stainless-steel component made up of two parts – the NN bar itself and an adapter – has been assembled and lifted to its permanent position in the area surrounding the Target Monolith vessel. 39 port tubes, providing the housing for the first few metres of the neutron extraction system, are subject to welding to the Monolith vessel.
About 50 penetrations in the 1,3-meter-thick heavy concrete wall separating the Technical Gallery from the Active Cells, are being closed with steel plugs. In-kind partner UKAEA, responsible for the remote handling system for the Active Cells, are performing the works.
The cladding for the cantilever roof encapsulating the Target building and adjacent instrument halls, is close to completion. Just a few gaps and joints remain to be equipped with panels of aluminium lamellae.
LOKI and FREIA
This week the installation of a common control hutch building for the instruments LOKI and FREIA started inside Experimental Hall 2. Built up of precast concrete blocks, the hutches will serve as sample preparation area and control room for the two instruments. LOKI and FREIA are both in-kind contributions from STFC, UK.
Installation of base components for the LOKI beamline and cave continues. A frame of steel plates is in place for the cave perimeter and concrete plinths have been placed close to the bunker. Drilling for floor anchors continue, as more base plates and the structure for the detector vessel to be placed inside the cave are added.
A staircase to roof level has been installed for the BIFROST instrument cave inside Experimental Hall 3. The spectrometer vessel, which has been lifted out from the cave, will be subject to cadmium coating.
Shielding blocks are stored inside the fan shaped E02 beamline gallery, with space for eight beamlines to the instruments in Experimental Hall 3. For CSPEC and BIFROST, parts of the lower shielding blocks are in place. Upper shielding blocks have started for the CSPEC beamline.
All green landscaping for the 70 hectares ESS site is complete, including planting and road signs, and next week the sectional handover inspection for remaining civil areas is planned.
Two thirds of the site consist of green landscaped areas.